OpenStudy (anonymous):

Inverse laplace of 1/{[(s-1)^2] + 1}?

OpenStudy (therealmeeeee):

Partial fractions: 1/[s(s+1)^2] = A/s + B/(s+1) + C/(s+1)^2 1 = A(s+1)^2 + Bs(s+1) + Cs 1 = As^2 + 2As + A + Bs^2 + Bs + Cs Equate like powers of s: 1 = A 0 = 2A + B + C 0 = A + B Together these imply that B = -1 and C = -1. 1/[s(s+1)^2] = 1/s - 1/(s+1) - 1/(s+1)^2 Now go to a Laplace transform table, and you get f(t) = 1 - e^(-t) - te^(-t)

OpenStudy (therealmeeeee):

do this help

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

yeah it helps but the equation i wrote is totally different.. I solved it inverse laplace of 1/[(s-1)^2 +1]=e^tcost any way thanks alot ! :)

OpenStudy (anonymous):

=( e^t ) cos (t) **

OpenStudy (therealmeeeee):

now click best response